# Orbit lengths for promotion on 2-row and near-hook tableaux

### v1

@@ 1 Introduction
Promotion is an operation on standard Young tableaux introduced by Schützeberger in [Sch72] that is closely related to jeu-de-taquin and has been extensively studied for rectangular tableax. In [Hai92], Haiman shows that for rectangular and staircase tableaux, the order of promotion (i.e. the number of times it must be repeated to recover the starting tableau) always divides the total number of boxes. Rhoades [Rho10] shows using representation theory that the set of orbit lengths for promotion of rectangular tableaux exhibits the cyclic sieving phenomenon (CSP) of Reiner, Stanton, and White [RSW04]. Alternative representation theoretic proofs for that CSP are given by Westbury [Wes16, Wes19], and geometric proofs are given by Fontaine and Kamnitzer [FK14] and Purbhoo [Pur13]. Schank, Sekheri, and Djermane [SSD14] prove a similar CSP theorem for staircase tableaux. A visual interpretation for the rectangle-case CSP in terms of webs is given by Petersen, Pylyavskyy, and Rhoades [PPR09] for the 2-row and 3-row cases, and by Gaetz, Pechenik, Pfannerer, Striker, and Swanson [GPPSS23, GPPSS24] for the 4-row case. Catania, Kendrick, Russell, and Tymoczko [CKRT25] give an algorithm for determining the exact orbit length for promotion given a specific rectangular tableau.
However, the orbit lengths for promotion on tableaux of other shapes are more complicated and have generally been believed not to have a nice structure, or an associated CSP. For instance, Striker [Str17] gives as an example that the order of promotion for tableaux of the 2-row shape (8,6) is 7,554,844,752. However, we show here that the orbit lengths for 2-row tableaux and certain other non-rectangular tableaux do in fact have an understandable structure which we describe precisely, and we also show that there is in fact a CSP closely related to the CSP for rectangular tableaux if one restricts to certain subsets of 2-row tableaux. The large-looking promotion order 7,554,844,752 from that example should really be thought of as the least common multiple of a number of much smaller orbit lengths for various subsets of tableaux, which we will describe precisely in §5.
@@ 1.1 Main results and organization
  • In §4, we prove a CSP theorem (Theorem 4.1.2) on the subsets of 2-row tableaux of a given shape such that all runs of consecutive numbers in the bottom row are the same length (with some subtleties about runs being allowed to “wrap around” from $n$ back to the beginning, and about the runs needing to be sufficiently far apart from each other). We shwo in particular that the orbit lengths divide certain monic linear polynomials in $n$. Then in Theorem 4.2.2, we show that the CSP polynomial almost matches the major index generating function for the tableaux in these subsets.
@@ 1.2 Future directions
There are many more questions to ask here. Based on experiments in Sage [Ste+25], the orbit lengths for the tableaux $T[n]$ built from a fixed starting tableau $T$ are not always divisors of fixed polynomials in $n$ (although they are for the cases studied here). However, it appears the orbit lengths may always be divisors of fixed quasipolynomials in $n$, meaning piecewise functions of $n$ where the cases depend only on the value of $n$ mod some fixed number, and where the function is a polynomial in each case. We pose this as a conjecture:
@@
Our work here shows that in the 2-row case, the near-hook case, and the “generic case” for arbitrary tableau shapes, the orbit lengths are in fact always divisors of certain polynomials rather than quasipolynomials, but it would be interesting to know when exactly this is the case for tableaux of other shapes:
@@
For which tableaux $T$ does the orbit length for $T[n]$ always divide a fixed polynomial in $n$? Are there shapes aside from 2-row and near-hook tableaux for which the orbit length always divides a fixed polynomial rather than a fixed quasipolynomial for all tableaux of that shape?
The overarching question would then be to describe how to compute these polynomials and/or quasipolynomials in general:
Question 1.2.3.
@@ 2.1 Young tableaux and promotion
A partition $\lambda=\lambda_{1}\dots\lambda_{\ell}$ is a nondecreasing sequence of positive integers $\lambda_{1}\geq\dots\geq\lambda_{\ell}.$ We call $\ell=\ell(\lambda)$ is its length, and $|\lambda|:=\lambda_{1}+\dots+\lambda_{\ell}$ is its size. For $n>|\lambda|+\lambda_{1}$, we write $\lambda[n]$ for the partition obtained from $\lambda$ by appending a new largest part of size $n-|\lambda|$ to get a partition of $n$.
@@ 3.1 Generic case orbit lengths
$\frac{\textnormal{lcm}(|T|,\textnormal{pr}(T))}{|T|}\cdot(n-1).$
@@ 11 $\cdots$ $\geq 3$ $\cdots$ $\vdots$ $\vdots$
total promotion steps, we will have completed $\textnormal{lcm}(|T|,\textnormal{pr}(T))/|T|$ of these rounds that each include $|T|$ steps satisfying (1), for a total of $\textnormal{lcm}(|T|,\textnormal{pr}(T))$ steps satisfying (1). Since the total number of promotion steps will is a multiple of n1n-1, (1) will be satisfied. (2) will also be satisfied because the number of steps so far satisfying (1) will a multiple of $|T|$, and (3) will hold because the number of steps satisfying (1) will be a multiple of $\textnormal{pr}(T)$. Since all three conditions hold, we will be back at the starting tableau $T[n]$ after $\textnormal{lcm}(|T|,\textnormal{pr}(T))/|T|\cdot(n-1)$ steps, so that number of steps must be a multiple of $\textnormal{pr}(T[n])$. ∎
@@ 11 $\cdots$ $k+1$ $\cdots$
Since we know that $\textnormal{{T}}(n,\ell,r)$ is closed under promotion and that $\textnormal{pr}(T[n])\mid(n-\ell)$ for every $T[n]\in\textnormal{{T}}(n,\ell,r)$, promotion gives an action of the cyclic group $\mathbb{Z}_{n-\ell}=\langle\mathcal{P}\rangle$ on $\textnormal{{T}}(n,\ell,r)$, so it remains to count the number of fixed points of of $\mathcal{P}^{(n-\ell)/d}$ for each $d\mid(n-\ell)$, and to check that it matches the value we get from plugging in a primitive $d^{\text{th}}$ root of unity to the claimed CSP polynomial.
@@ 5.4 Recovering a tableau from its set of tracks
Compare some orbit $\mathcal{O}$ within $\textnormal{{T}}(n,\vec{\ell},\vec{r})$ to another orbit $\mathcal{O}^{\prime}$ where the $i^{\text{th}}$ track starts shifted by some amount $s_{i}\leq\ell_{d}$ in $\mathcal{O}^{\prime}$ compared to where it starts in $\mathcal{O}$, while all other tracks start in the same positions in $\mathcal{O}$ and $\mathcal{O^{\prime}}$. We claim that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ have the length. To see this, we will show that throughout $\mathcal{O}$ and $\mathcal{O}^{\prime}$, the $i^{\text{th}}$ track will always stay offset by the same amount in $\mathcal{O}^{\prime}$ relative to $\mathcal{O}$, while
@@
Now, given two orbits within the size $P_{d}(n,\vec{\ell},\vec{r}$) subset of $\textnormal{{T}}(n,\vec{\ell},\vec{r})$ that corresponds to the a particular set of gap sequences, we can show that those orbits always have the same length by turning one orbit into the other through a series of these steps of shifting one track at a time by one position while keeping the other tracks fixed. If we run into a bottleneck issue when trying to shift one of the tracks by one, then we can just wait some number of promotion steps until the rainbow that would cause the bottleneck issue has finished crossing the boundary, and then shift the track by one at that position in the orbit instead to show that the orbits have the same length. It follows that $\textnormal{{T}}(n,\vec{\ell},\vec{r})$ can be partitioned into subsets of size $P_{d}(n,\vec{\ell},\vec{r})$, each of which is closed under promotion and can be partitioned into promotion orbits that all have the same length. Thus, this common orbit length must be a divisor of $P_{d}(n,\vec{\ell},\vec{r}).$
@@
matching what we get from plugging in $q=1$ to the CSP polynomial in Theorems 4.1.2 and 4.2.2.
@@ 6.1 Linear orbit lengths for near-hooks
If we start from the beginning of the first run, adding $\ell_{1}-1\pmod{n-1}$ will take us to the end of the first run. Then adding $g_{2}\pmod{n-1}$ will take us to the start of the second run, adding $\ell_{2}-1\pmod{n-1}$ will take us to the end of the second run, and so on. Adding $\sum_{i=1}^{r}(\ell_{i}-1+r_{i})\pmod{n-1}$ almost takes us back to the same number from the first column that we started at, so it would seem that the sum $\sum_{i=1}^{r}(\ell_{i}-1+r_{i})$ should equal n1n-1, but actually it equals $n$ because the extra number in the top right box of $T$ gets double counted in both one of the $\ell_{i}-1$ terms and in one of the $g_{i}$ terms, since the gap containing the top right number is counted as the difference between the number before the top right number in the first column and the number after the top right number in the first column, rather than between the top right number itself and either the number before or after it. Thus, we have $\sum_{i=1}^{r}(\ell_{i}-1+g_{i})=n+r,$ so $\sum_{i=1}^{r}(\ell_{i}+g_{i})=n+r.$
@@ 11 $\cdots$ 33 $\ell_{1}+2$ 44 $\vdots$ $\ell_{1}+1$ $\vdots$
Next, all the entries belonging to other runs in $T$ will simply slide up and wrap around from being a 2 at the top of $T$ to being an $n$ at the bottom of $T$, until we get to a point where the portion of the first run that is in the first column starts just below the top left box of $T$. Then the smallest entry of the first run will slide over to the left to become the top left entry of $T$, so the first run will be entirely in the first column of $T$. The first entry of the first run slides left at the step where the last entry of the $r^{\text{th}}$ run is a 2 and then wraps around to become an $n$ in the top right corner of $T$. Since the mod n1n-1 gap between the last entry of the $r^{\text{th}}$ run within the left column and the first entry of the first run within the left column must be $g_{1}$, the first entry of the first run within the column is $g_{1}+2$ at that step, so the entry in the top right box is $g_{1}+1$, and then is $g_{1}$ when it slides left. In the figure below, entries of the first run are still highlighted in green, while entries of the $r^{\text{th}}$ run are highlighted in yellow:

### v2

@@ 1 Introduction
Promotion is an operation on standard Young tableaux introduced by Schützeberger in [11] that is closely related to jeu-de-taquin and has been extensively studied for rectangular tableaux. In [5], Haiman shows that for rectangular and staircase tableaux, the order of promotion (i.e. the number of times it must be repeated to recover the starting tableau) always divides the total number of boxes. Rhoades [9] shows using representation theory that the set of orbit lengths for promotion of rectangular tableaux exhibits the cyclic sieving phenomenon (CSP) of Reiner, Stanton, and White [8]. Alternative representation theoretic proofs for that CSP are given by Westbury [14, 15], and geometric proofs are given by Fontaine and Kamnitzer [2] and Purbhoo [7]. Schank, Sekheri, and Djermane [12] prove a similar CSP theorem for staircase tableaux. A visual interpretation for the rectangle-case CSP in terms of webs is given by Petersen, Pylyavskyy, and Rhoades [6] for the 2-row and 3-row cases, and by Gaetz, Pechenik, Pfannerer, Striker, and Swanson [3, 4] for the 4-row case. Catania, Kendrick, Russell, and Tymoczko [1] give an algorithm for determining the exact orbit length for promotion given a specific rectangular tableau.
However, the orbit lengths for promotion on tableaux of other shapes are more complicated and have generally been believed not to have a nice structure, or an associated CSP. For instance, Striker [13] gives as an example that the order of promotion for tableaux of the 2-row shape (8,6) is 7,554,844,752. However, we show here that the orbit lengths for 2-row tableaux and certain other non-rectangular tableaux do in fact have an understandable structure which we describe precisely, and we also show that there is in fact a CSP closely related to the CSP for rectangular tableaux if one restricts to certain subsets of 2-row tableaux. The large-looking promotion order 7,554,844,752 from that example should really be thought of as the least common multiple of a number of much smaller orbit lengths for various subsets of tableaux, which we will describe precisely in §5.
@@ 1.1 Main results and organization
  • In §4, we prove a CSP theorem (Theorem 4.1.2) on the subsets of 2-row tableaux of a given shape such that all runs of consecutive numbers in the bottom row are the same length (with some subtleties about runs being allowed to “wrap around” from $n$ back to the beginning, and about the runs needing to be sufficiently far apart from each other). We show in particular that the orbit lengths divide certain monic linear polynomials in $n$. Then in Theorem 4.2.3, we show that the CSP polynomial almost matches the major index generating function for the tableaux in these subsets.
@@ 1.2 Future directions
There are many more questions to ask here. Based on experiments in Sage [10], the orbit lengths for the tableaux $T[n]$ built from a fixed starting tableau $T$ are best described not as divisors of fixed polynomials in $n$ (although they are for the cases studied here), but rather as divisors of fixed quasipolynomials in $n$, meaning piecewise functions of $n$ where the cases depend only on the value of $n$ mod some fixed number, and where the function is a polynomial in each case. We pose this as a conjecture:
@@
One could of course still say in such a case that the orbit length is a divisor of a fixed polynomial obtained by multiplying all the cases of the quasipolynomial, but it is more informative to describe the orbit lengths in terms of the quasipolynomial. The overarching question would then be to describe how to compute these polynomials and/or quasipolynomials in general:
@@ 2.1 Young tableaux and promotion
A partition $\lambda=\lambda_{1}\dots\lambda_{\ell}$ is a nonincreasing sequence of positive integers $\lambda_{1}\geq\dots\geq\lambda_{\ell}.$ We call $\ell=\ell(\lambda)$ is its length, and $|\lambda|:=\lambda_{1}+\dots+\lambda_{\ell}$ is its size. For $n>|\lambda|+\lambda_{1}$, we write $\lambda[n]$ for the partition obtained from $\lambda$ by appending a new largest part of size $n-|\lambda|$ to get a partition of $n$.
@@ 3.1 Generic case orbit lengths
$\frac{\textnormal{lcm}(|T|,\textnormal{pr}(\textnormal{{SYT}}(T)))}{|T|}\cdot(n-1).$
@@ 11 $\cdots$ $\geq 3$ $\cdots$ $\vdots$ $\vdots$
total promotion steps, we will have completed $\textnormal{lcm}(|T|,\textnormal{pr}(T))/|T|$ of these rounds that each include $|T|$ steps satisfying (1), for a total of $\textnormal{lcm}(|T|,\textnormal{pr}(T))$ steps satisfying (1). Since the total number of promotion steps will be a multiple of n1n-1, (1) will be satisfied. (2) will also be satisfied because the number of steps so far satisfying (1) will be a multiple of $|T|$, and (3) will hold because the number of steps satisfying (1) will be a multiple of $\textnormal{pr}(T)$. Since all three conditions hold, we will be back at the starting tableau $T[n]$ after $\textnormal{lcm}(|T|,\textnormal{pr}(T))/|T|\cdot(n-1)$ steps, so that number of steps must be a multiple of $\textnormal{pr}(T[n])$. ∎
Example 3.1.3.
Let $T=$ 44 66 88      and $n=8,$ so $\textnormal{{SYT}}(T)=$ 11 22 33      and
T[8]=
1 7
4 6
8
[…]
Then $\mathcal{P}(\textnormal{{SYT}}(T))=$ 11 33 22      and $\mathcal{P}(\mathcal{P}(\textnormal{{SYT}}(T))$ = 11 22 33      $\textnormal{{SYT}}(T)$, so $\textnormal{pr}(\textnormal{{SYT}}(T))=2$. So we should have
$\textnormal{pr}(T[8])\mid\frac{\textnormal{lcm}(3,2)}{3}\cdot(8-1)=2\cdot 7=14.$
If we track the entry that starts as the 8 (shown in orange), it will go through 2 full cycles during the orbit, going from 8 down to 2. During each of those cycles, there are 3 steps (shown in yellow) where an entry of $T$ becomes 2 and $\textnormal{{SYT}}(T)$ changes in the next step. Thus, $\textnormal{{SYT}}(T)$ goes through 6 promotion steps over the orbit, so it ends up back at its starting state since $\textnormal{pr}(\textnormal{{SYT}}(T))=2$.
11 77 44 66 88 11 22 33 $\longrightarrow$ 11 88 33 55 77 11 22 33 $\longrightarrow$ 11 88 22 44 66 11 22 33
$\longrightarrow$ 11 77 33 88 55 11 33 22 $\longrightarrow$ 11 88 22 77 44 11 33 22 $\longrightarrow$ 11 77 33 66 88 11 22 33
$\longrightarrow$ 11 88 22 55 77 11 22 33
$\longrightarrow$ 11 77 44 88 66 11 33 22 $\longrightarrow$ 11 88 33 77 55 11 33 22 $\longrightarrow$ 11 88 22 66 44 11 33 22
$\longrightarrow$ 11 77 33 55 88 11 22 33 $\longrightarrow$ 11 88 22 44 77 11 22 33 $\longrightarrow$ 11 77 33 88 66 11 33 22
$\longrightarrow$ 11 88 22 77 55 11 33 22
@@ 11 $\cdots$ $k+1$ $\cdots$
Since we know that $\textnormal{{T}}(n,\ell,r)$ is closed under promotion and that $\textnormal{pr}(T[n])\mid(n-\ell)$ for every $T[n]\in\textnormal{{T}}(n,\ell,r)$, promotion gives an action of the cyclic group $\mathbb{Z}_{n-\ell}=\langle\mathcal{P}\rangle$ on $\textnormal{{T}}(n,\ell,r)$, so it remains to count the number of fixed points of $\mathcal{P}^{(n-\ell)/d}$ for each $d\mid(n-\ell)$, and to check that it matches the value we get from plugging in a primitive $d^{\text{th}}$ root of unity to the claimed CSP polynomial.
@@
Example 4.1.6.
Below is an orbit we get for $\ell=3$ and $n=16$, with one run shown in orange and the other in green. The orbit length is $n-\ell=16-3=13.$ For each run, the “first” number cycles through the 13 values $16,15,\dots,4.$
11 $16$ 44 $12$ $\longrightarrow$ 11 $15$ 33 $16$
$\longrightarrow$ 11 $14$ 22 $16$ $\longrightarrow$ 11 $13$ 77 $16$
$\longrightarrow$ 11 $16$ 66 $15$ $\longrightarrow$ 11 $16$ 55 $14$
$\longrightarrow$ 11 $16$ 44 $13$ $\longrightarrow$ 11 $15$ 33 $16$
$\longrightarrow$ 11 $14$ 22 $16$ $\longrightarrow$ 11 $13$ 88 $16$
$\longrightarrow$ 11 $16$ 77 $15$ $\longrightarrow$ 11 $16$ 66 $14$
$\longrightarrow$ 11 $16$ 55 $13$
@@ 4.2 Major index interpretation of the cyclic sieving polynomial
Example 4.2.2.
Let $T[15]$ be the second tableau from Example 4.1.6:
11 $15$ 33 $16$
To compute $\textnormal{maj}(T[15])$, we count 2, 8, and 15 as all being descents, but to compute $\textnormal{maj}_{3}(T[15])$, we only count 8 and 15 as descents, since $2<3$. Thus, $\textnormal{maj}(T[15])=2+8+15=25$ while $\textnormal{maj}_{3}(T[15])=8+15=23.$
@@ 5.1 Determining the run lengths from a tableau
Example 5.1.1.
For the tableau $T[22]$ below, the gaps between runs are large enough that the runs are just the sequences of consecutive numbers in the bottom row:
11 $22$ 44 $18$
The longest run is $(4,5,6)$ of length 3, so $\ell_{1}=3$ and $r_{1}=1$, the next longest runs are $(13,14)$ and $(17,18)$, each of length 2, so $\ell_{2}=2$ and $r_{2}=2$, and the final run is $(10)$ of length 1, so $\ell_{3}=r_{3}=1$. Thus, $\vec{\ell}=(3,2,1)$ and $\vec{r}=(1,2,1).$
@@ 5.4 Recovering a tableau from its set of tracks
Compare some orbit $\mathcal{O}$ within $\textnormal{{T}}(n,\vec{\ell},\vec{r})$ to another orbit $\mathcal{O}^{\prime}$ where the $i^{\text{th}}$ track starts shifted by some amount $s_{i}\leq\ell_{d}$ in $\mathcal{O}^{\prime}$ compared to where it starts in $\mathcal{O}$, while all other tracks start in the same positions in $\mathcal{O}$ and $\mathcal{O^{\prime}}$. We claim that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ have the same length. To see this, we will show that throughout $\mathcal{O}$ and $\mathcal{O}^{\prime}$, the $i^{\text{th}}$ track will always stay offset by the same amount in $\mathcal{O}^{\prime}$ relative to $\mathcal{O}$, while
@@
Now, given two orbits within the size $P_{d}(n,\vec{\ell},\vec{r}$) subset of $\textnormal{{T}}(n,\vec{\ell},\vec{r})$ that corresponds to a particular set of gap sequences, we can show that those orbits always have the same length by turning one orbit into the other through a series of these steps of shifting one track at a time by one position while keeping the other tracks fixed. If we run into a bottleneck issue when trying to shift one of the tracks by one, then we can just wait some number of promotion steps until the rainbow that would cause the bottleneck issue has finished crossing the boundary, and then shift the track by one at that position in the orbit instead to show that the orbits have the same length. It follows that $\textnormal{{T}}(n,\vec{\ell},\vec{r})$ can be partitioned into subsets of size $P_{d}(n,\vec{\ell},\vec{r})$, each of which is closed under promotion and can be partitioned into promotion orbits that all have the same length. Thus, this common orbit length must be a divisor of $P_{d}(n,\vec{\ell},\vec{r}).$
@@
matching what we get from plugging in $q=1$ to the CSP polynomial in Theorems 4.1.2 and 4.2.3.
@@ 6.1 Linear orbit lengths for near-hooks
If we start from the beginning of the first run, adding $\ell_{1}-1\pmod{n-1}$ will take us to the end of the first run. Then adding $g_{2}\pmod{n-1}$ will take us to the start of the second run, adding $\ell_{2}-1\pmod{n-1}$ will take us to the end of the second run, and so on. Adding $\sum_{i=1}^{r}(\ell_{i}-1+g_{i})\pmod{n-1}$ almost takes us back to the same number from the first column that we started at, so it would seem that the sum $\sum_{i=1}^{r}(\ell_{i}-1+g_{i})$ should equal n1n-1, but actually it equals $n$ because the extra number in the top right box of $T$ gets double counted in both one of the $\ell_{i}-1$ terms and in one of the $g_{i}$ terms, since the gap containing the top right number is counted as the difference between the number before the top right number in the first column and the number after the top right number in the first column, rather than between the top right number itself and either the number before or after it. Thus, we have $\sum_{i=1}^{r}(\ell_{i}-1+g_{i})=n+r,$ so $\sum_{i=1}^{r}(\ell_{i}+g_{i})=n+r.$
@@ 11 $\cdots$ 33 $\ell_{1}+2$ 44 $\vdots$ $\ell_{1}+1$ $\vdots$
Next, all the entries belonging to other runs in $T$ will simply slide up and wrap around from being a 2 at the top of $T$ to being an $n$ at the bottom of $T$, until we get to a point where the portion of the first run that is in the first column starts just below the top left box of $T$. Then the smallest entry of the first run will slide over to the left to become the top left entry of $T$, so the first run will be entirely in the first column of $T$. The first entry of the first run slides left at the step where the last entry of the $r^{\text{th}}$ run is a 2 and then wraps around to become an $n$ in the top right corner of $T$. Since the mod n1n-1 gap between the last entry of the $r^{\text{th}}$ run within the left column and the first entry of the first run within the left column must be $g_{1}$, the first entry of the first run within the column is $g_{1}+2$ at that step, so the entry in the top right box is $g_{1}+1$, and then is $g_{1}$ when it slides left. In the figure below, entries of the first run are still highlighted in green, while entries of the $r^{\text{th}}$ run are highlighted in orange:
@@ 11 $\cdots$ $\scriptstyle{n-\ell_{1}+1}$ $\vdots$ $\scriptstyle{n-\ell_{1}+2}$ $\vdots$ $n$
Example 6.1.6.
For the tableau $T[9]$ shown below, we have 2 runs, of lengths 2 and 3, shown in green and orange. At first the green run has its largest entry in the top right, then it has its smallest entry in the top right, then the orange run has its largest entry in the top right, then its smallest. Since $r=2$, we get an orbit length of $3\cdot 9-4=23.$
11 66 33 99 44 77 88 $\longrightarrow$ 11 99 22 88 33 66 77 $\longrightarrow$ 11 88 22 77 55 66 99 $\longrightarrow$ 11 77 44 66 55 88 99 $\longrightarrow$ 11 99 33 55 44 77 88
$\longrightarrow$ 11 88 22 99 33 66 77 $\longrightarrow$ 11 77 22 88 55 66 99 $\longrightarrow$ 11 77 44 77 55 88 99 $\longrightarrow$ 11 99 33 66 44 77 88 $\longrightarrow$ 11 99 22 55 33 66 77
$\longrightarrow$ 11 88 22 44 55 66 99 $\longrightarrow$ 11 77 33 99 44 55 88 $\longrightarrow$ 11 99 22 88 33 44 77 $\longrightarrow$ 11 88 22 77 33 66 99 $\longrightarrow$ 11 88 22 66 55 88 99
$\longrightarrow$ 11 66 44 55 77 88 99 $\longrightarrow$ 11 99 33 44 66 77 88 $\longrightarrow$ 11 88 22 99 55 66 77 $\longrightarrow$ 11 88 44 88 55 66 99 $\longrightarrow$ 11 99 33 77 44 55 88
$\longrightarrow$ 11 99 22 66 33 44 77 $\longrightarrow$ 11 99 22 55 33 66 99 $\longrightarrow$ 11 77 22 44 55 88 99